You are lying 120 ft away from a tree that is 50 feet tall. You look up at the top of the tree. Approximately how far is your hear from the top of the tree in a straight line?

please help me understand how to solve this problem. Kevin.

Answer by richwmiller(17219)


what is the hypotenuse of the right triangle with legs 50 ft and 120 feet?
a^2+b^2=c^2
50^2+120^2=c^2
But you should see that this is a 10 times multiple of the classic 5,12,13 right triangle.
So the hypotenuse is 130 feet.